OpenStudy (angelwings996):

I have started this problem myself but as I was working I got stuck and can't move on. Use the quadratic formula to solve the equation x^2 - 7x - 6 = 0

OpenStudy (anonymous):

already teach u

OpenStudy (angelwings996):

What ?

OpenStudy (anonymous):

\[x^2-7x-6\]

OpenStudy (angelwings996):

Yes but this is a different kind of problem. Do you want to see what I have done so far ?

OpenStudy (anonymous):

\[a=1\rightarrow b=-7 \rightarrow c=-6\]

OpenStudy (angelwings996):

Yes I already plugged them into the formula

OpenStudy (hba):

Show us your working then :)

OpenStudy (anonymous):

wht hv u got?

OpenStudy (hba):

Better use latex :)

OpenStudy (angelwings996):

Okay, so the quadratic formula is \[x = -b \pm \sqrt{b ^{2} - 4ac} / 2a\]

OpenStudy (anonymous):

yap

OpenStudy (hba):

The 2a is divided by the whole thing.

OpenStudy (angelwings996):

I plugged a, b, and c in and got \[x = -(-7) \pm \sqrt{-7^{2} - 4(1)(-6)} / 2(1)\]

OpenStudy (angelwings996):

Yes I know that the 2a is multiplied by the whole thing

OpenStudy (anonymous):

ur r8

OpenStudy (hba):

Okay :) But it is divided not multiplied :P

OpenStudy (angelwings996):

Then I got \[x = 7 \pm \sqrt{49 + 24} / 2\]

OpenStudy (angelwings996):

Sorry I meant divided by lol

OpenStudy (hba):

Okay good going :)

OpenStudy (anonymous):

good

OpenStudy (angelwings996):

Then I got \[x = 7 \pm \sqrt{73} / 2\] and the sqaure root of 73 is 8.544

OpenStudy (angelwings996):

Then I got stuck because I don't know what to do now

OpenStudy (angelwings996):

would I do 7 + 8.544 then divide the whole thing by 2 then do the same thing but subtract instead of add ?

OpenStudy (hba):

Well now you can distribute it